OpenStudy (anonymous):

Question, How do you find the equation of a line when given the x and y intercept?

OpenStudy (mertsj):

Use the two points to find the slope. You are given the y intercept. Now you know the slope and the y intercept. Use slope-intercept form and write the equation.

OpenStudy (anonymous):

I just plug in the slope for M?

OpenStudy (mertsj):

yes

OpenStudy (anonymous):

So if I have x intercept -5 and y intecept -2 the equation would look like this? y = -2/5x - 2

OpenStudy (tkhunny):

Those familiar with the "Intercept Form" will solve this a LOT faster. "x intercept -5 and y intecept -2" \(\dfrac{x}{-5} + \dfrac{y}{-2} = 1\) -- Done! If you need some other form after that, well, algebra away!